18x - 12y = 18 x = y - 12 what is x? What is y?
Answer by checkley77(12844)

18x - 12y = 18
x = y - 12
Substituting (y-12) for x in the first equation we get:
18(y-12)-12y=18
18y-216-12y=18
6y=18+216
6y=234
y=234/6
y=39 ans.
x=39-12
x=27 ans.
Proof:
18*27-12*39=18
486-468=18
18=18